40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Fri Dec 20 21:58:14 2019, Koji, Update, PD QE, PD TEC driver / A2P6 aligned / Lens moved P_20191220_192440_vHDR_On.jpgP_20191220_180929_vHDR_On.jpg
    Reply  Fri Dec 20 22:26:19 2019, Koji, Update, PD QE, PD TEC driver / A2P6 aligned / Lens moved 6x
Message ID: 2483     Entry time: Fri Dec 20 22:26:19 2019     In reply to: 2482
Author: Koji 
Type: Update 
Category: PD QE 
Subject: PD TEC driver / A2P6 aligned / Lens moved 

The QEs were measured at 293K, 239K, 232K, and 293K again. The cooling was provided by the PD TEC.  At each temperature, the incident power was changed from 30uW to 1mW to see the dependence of the QE on the incident power to check the possible saturation.

The QE was 79~81% (the window T=96.6% was already compensated). I'm not 100% sure this 1% variation in the plateau is real or due to insufficient calibration of the REF PD.
The REF PD was calibrated at 1mW at 100mA injection current to the laser.

No obvious saturation was observed.

We can cool the PD with LN2 and we should make a careful alignment of the beam at each temperature.

Attachment 1: Sb3513_A2P6_DarkCurrent_293K.pdf  130 kB  | Hide | Hide all | Show all
Sb3513_A2P6_DarkCurrent_293K.pdf
Attachment 2: Sb3513_A2P6_DarkCurrent_239K.pdf  124 kB  | Show | Hide all | Show all
Attachment 3: Sb3513_A2P6_DarkCurrent_232K.pdf  124 kB  | Hide | Hide all | Show all
Sb3513_A2P6_DarkCurrent_232K.pdf
Attachment 4: Sb3513_A2P6_DarkCurrent_293K_2.pdf  124 kB  | Hide | Hide all | Show all
Sb3513_A2P6_DarkCurrent_293K_2.pdf
Attachment 5: Sb3513_A2P6_DarkCurrent_Comparison.pdf  119 kB  | Hide | Hide all | Show all
Sb3513_A2P6_DarkCurrent_Comparison.pdf
Attachment 6: 191220_3513A2P6.zip  1.573 MB
ELOG V3.1.3-